OpenStudy (anonymous):

a camp site is 9.41 miles from a point directly below the mountain top. if the angle of elevation is 12 degrees from the camp to the top of the mountain, how high is the mountain?

OpenStudy (cornzyblack):

Ok first uhmm we make use of SohCahToa Tan x = opposite/adjacent Tan 15 = y/9.41 y = tan 15 * 9.41 y = 2.52 miles So the mountain is 2.52 miles high